• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 21:47:31 on November 18, 2008

    Entry number 249628

    This entry is a followup to: 249538

    keyword=BB Cerenkov ADC<->TDC mapping problem fixed

    There was indeed a partial scrambling of the signals running to the TDC discriminators. It's now fixed.

    The dark-current signal on PMT 17 was considerably lower rate than on the other PMTs... Amplitudes looked OK though. Not sure if it's dieing or what.

    A copy of this log entry has been emailed to: meziani